### v1

@@ 3.1 Survey demography
By default we simulate over the whole DESI footprint, however if a specific area is to be simulated, such as fractions of the DESI footprint or the footprint from other surveys, we use an input number density function tabulated into HEALpix pixels [60] including only those pixels that cover the footprint to be simulated. Quickquasars then uses this information and random samples the number of objects from the available skewers to match the observed number density. We used this sub-sampling method as a first approximation for EDR+M2 mocks only, even though in principle it could affect the Lyman-$\alpha$ – QSO cross and QSO auto-correlations at large scales for the HEALpix pixel size used. [^12] We use HEALpix pixel of nside=16 which corresponds to a $\approx 250$ Mpc/h scale at redshift z=2.3. This sub-sampling method differs from the one used for eBOSS mocks in the DR16 analysis validation [10] and therefore does not affect its results. A more suitable method will be studied for future releases of DESI mocks.
@@
We include two options to assign exposure times. If it is not expected that all the simulated quasars have the same exposure time, we use an exposure time probability distribution function computed for each HEALpix pixel of the footprint to randomly assign a multiple of 1000 seconds to the selected quasars. Otherwise, we assign all the quasars the same exposure time, 4000 seconds in the case of full DESI survey mocks.
@@ 3.2.3 Variance smoothing
This smoothing step is not part of specsim, so we emulate this process in quickquasars as follows: We subtract the source electron shot noise contribution from the total variance estimated by specsim, explained in section 3.2.2. We smooth this source electron contribution with a Gaussian kernel using Fast Fourier Transforms (FFTs), where the input array is padded with boundary values to prevent periodicity from distorting values at the edges. Finally, this smoothed source contribution is added to the total variance. The standard deviation of the Gaussian kernel $\sigma$ is an input parameter in quickquasars, and based on the spectroscopic data processing pipeline it is set to $10~{}\text{\AA}$ by default.
@@ 3.2.4 Observing conditions
The atmosphere model presented in section 3.2.1 depends on the assumed observing conditions such as seeing, air mass, moon illumination fraction, moon altitude and moon separation from the tile being simulated. For Lyman-$\alpha$ mocks we use the DARK program setup where the seeing is set to $1.1\ \rm{arcsec}$, air mass to 1.0, and we set the moon illumination fraction to 0 while the moon altitude and separation from the tile are set to $-60\ \rm{degrees}$ and $180\ \rm{degrees}$, respectively. While these conditions might not be particularly realistic, they correspond to how the effective exposure time is defined for DESI observations. In other words, a 1000 seconds exposure simulated at these conditions provides the expected signal to noise ratio that an actual observation under different conditions aims to achieve.
@@ 3.2.5 The resolution matrix
The DESI spectrograph resolution is encoded in our data products in the form of a "resolution matrix" that is saved in the output spectra files. In observed data this matrix is a result of the spectroscopic extraction algorithm and encodes the dispersion along wavelength due to the spectrograph finite PSF and the CCD pixel size. More details can be found in section 4.5 of [63]. This band-diagonal matrix  [^15] The resolution matrix elements far from the diagonal are zero since it combines only flux values from neighboring wavelength. applied to a high resolution spectrum results in a spectrum at the resolution of the observations. For mocks, each row of the resolution matrix can be approximated as a Gaussian of parameter $\sigma=(\sigma_{LSF}^{2}+\Delta\lambda^{2}/12)^{1/2}$, where $\sigma_{LSF}=0.73$Å at 4000Å and $\Delta\lambda=0.8$Å is the output wavelength bin size.
@@ 4 DESI EDR+M2 Lyman-$\alpha$ mocks comparison with data
In this section we present the DESI Lyman-$\alpha$ EDR+M2 mocks, a collection of 40 different mock datasets: 20 for LyaCoLoRe and 20 for Saclay raw mocks, of which 10 mocks contain only Lyman-$\alpha$ absorption and 10 additionally include HCDs, BALs and metals. The metals were extracted directly from the transmission files for the LyaCoLoRe mocks and added by quickquasars for the Saclay ones. We include absorption due to Lyman-$\beta$, Si II(1190), Si II(1193), Si II(1260) and Si III(1207).
To produce all of these mocks we extracted the observed footprint, quasar number density and exposure time distribution from the same DESI quasar catalog used for the observed data 3D correlation functions measurement [64]. This catalog combines the data of DESI Early Data Release (EDR), collected during the survey validation phase, and the first two months of the Main Survey (DESI-M2). The resulting observed data sample consists of 318k QSO targets within the $60<z<$ 6 redshift range. Next, we show a comparison of the results of these mocks to those of the observed data.
@@ 4.1 Demographics
For the mocks we only consider QSOs with redshift $3.81.8<z<$ 3.8, which results in a sample of 141k targets including 107k Lyman-$\alpha$ QSOs at $z>2.1$. We ignore targets in the regions explored during the EDR phase that fall outside the main DESI footprint (far north and far south) since these regions are not considered by the raw mocks. This makes no impact on our comparison since these targets represent only 0.9% of the sample. The resulting footprint of the EDR+M2 dataset with the considerations previously mentioned is shown in figure 4. By construction all of the produced mocks follow the same footprint, QSO number density and redshift-magnitude distributions as the observed data with negligible variations between realizations.
By construction our mocks follow the redshift-magnitude distribution from figure 1 which is slightly different from that measured on the EDR+M2 sample. This is due the fact that the EDR sample includes fainter targets than those expected to be measured during the main DESI survey (e.g DESI-M2). [^16] During the survey validation phase an extension of the r-band magnitude limit was tested to study the redshift distribution and population of fainter objects. However, this does not affect the quality of our mocks since the amount of this type of target represents a small fraction of the sample as can be seen in the left panel of figure 5 at $M_{r}>23.1$.
@@
Figure 5 shows the observed EDR+M2 redshift (right) and r-band magnitude (left) distributions and the corresponding results of one mock realization of each LyaCoLoRe and Saclay mocks. The differences between mocks and data distributions are below the 1% level and we obtain negligible dispersion between each of the mock realizations. This validates the appropriateness of our methods of redshift sampling and random r-band magnitude assignation, described in section 3.1. We see a cutoff in the redshift distribution of the mocks at $z>3.8$ for LyaCoLoRe and $z>3.6$ for Saclay. This comes from the fact that we reach the redshift limit of the raw mocks at this point. The effect of these limits in our analysis and validations are negligible, since the QSOs at these redshifts represent around 1% of the observed sample.
To emulate the signal to noise ratio (SNR) of observed data we assign random exposure times to the simulated quasars by using a exposure time distribution function, as explained in section 3.1. The distribution function was computed from the effective exposure time of observed data defined by spectroscopic data processing pipeline by the equation Additionally, we use the measured throughput model achieved during the EDR phase (blue line in figure 3), which includes a dip feature at $\lambda\approx 440$ nm that affects the noise of the simulated spectra [62]. We also consider the effect of galactic extinction on spectra by adding an O’Donnell extinction model [65]. This in principle modifies the magnitude of the mock QSOs from the randomly assigned ones, and therefore modifies the resulting magnitude distribution. We address this issue by re-scaling all fluxes by a factor $F=f_{0}/f_{\rm{EBV}}$, where $f_{0}$ is the flux randomly assigned by our method, and $f_{\rm{EBV}}$ is the flux after galactic extinction has been applied.
Figure 6 shows the median SNR of a 75k QSO randomly selected sample in the Lyman-$\alpha$ forest region of observed data and a mock realization of both LyaCoLoRe and Saclay mocks. We have considered r-band magnitude and redshift bins of 0.5 and 0.1 width, respectively. We see a significant difference in the SNR of our mocks compared to data at bright low redshifts quasars ($z<2.2$, $M_{r}<22$), where we obtain higher SNR in our mocks than in observed data, which might be due to an underestimation in our instrumental noise model for bright low redshift quasars. We expect to study and improve on this issue for future mocks.
@@ 4.2 Correlation functions
The variance $\sigma_{q}$ for the auto-correlation has contributions from instrumental noise $\sigma_{\rm{pip}}$ and large-scale structure $\sigma_{\rm{LSS}}$
@@
where $\tilde{\sigma}_{\rm{pip},q}=\sigma_{\rm{pip},q}(\lambda)/\bar{F}C_{q}(\lambda)$, and $\eta(\lambda)$ is a correction factor found to be close to $1$ with small variation [66], allowing us to fix it to $1$ in our analysis. The parameter $\sigma_{\rm{LSS}}$ is fitted iteratively at the same time as the rest-frame continuum $\bar{C}(\lambda_{\rm{RF}})$ and the quasar spectral diversity parameters $a_{q}$ and $b_{q}$.
@@
Figure 7 shows the comparison of the distributions of the quasar spectral diversity parameters $a_{q}$ and $b_{q}/a_{q}$ as measured from data and mocks. We have restricted to those objects whose amplitude parameter is $a_{q}>0$. The difference between the $b_{q}/a_{q}$ distributions of data and mocks suggests that an update in the simqso broken power law model slopes and dispersions might be required to reflect better the quasar spectral diversity observed with DESI.
@@
where $\gamma_{\rm{Ly\alpha}}=2.9$ [68], $\gamma_{\rm{QSO}}=1.44$ [69] and $\sigma_{\rm{mod}}^{2}$ is an extra parameter introduced to modulate the contribution of the large-scale variance $\sigma_{\rm{LSS}}$. We have fixed the value $\sigma_{\rm{mod}}^{2}=7.5$ as this value was found to optimize the precision on the results of the correlation functions for the EDR+M2 dataset [66].
@@
Figure 9 shows the Lyman-$\alpha$ auto (top) and Lyman-$\alpha$ – QSO cross (bottom) correlation functions of the measurements presented by [64] in four ranges of $\mu=r_{\parallel}/r$. We also show individual correlations of the 10 mock realizations with contaminants for both LyaCoLoRe and Saclay mocks and their median value. We have multiplied the correlations by $r^{2}$ for better appreciation of the BAO peak. Overall the computed correlation functions of Saclay and LyaCoLoRe mocks are consistent with observations, however there are some clear differences between mocks and observed data, which we discuss next.
To begin with, we see significant differences at small scales specially at the $0.50<\mu<$ 0.5 range of the auto-correlation and the $0.80.5<\mu<$ 0.8 and $0.950.8<\mu<$ 0.95 ranges of the cross-correlation, which may be due to the difference between the Lyman-$\alpha$ and quasar biases of EDR+M2 observed data and the target values used to produce the raw mocks based on BOSS DR12 in the case of LyaCoLoRe and eBOSS DR16 in the case of Saclay. This will be studied in future DESI releases with higher signal to noise ratio. Some differences in the correlations may also be due to differences in DLA masking. As it was discussed in section 2.3, systems that go undetected by the DLA finder algorithm have a significant impact on the shape of the correlation functions, this is not the case for our mocks where we mask all of these features. Another possible contribution to the correlation differences may be the systematic errors on the measured redshift of observed quasars which were not introduced in our mocks. Systematic redshift errors alter the shape of the correlations functions in the $0.950.8<\mu<$ 0.95 and $10.95<\mu<$ 1 ranges for both auto and cross correlations [37]. The differences at the $1.00.95<\mu<$ 1.0 range at approximately 20 and 60 Mpc/h in the auto-correlation and 110 Mpc/h in the cross-correlation might be caused by the value chosen to include metals in our mocks, this will be further discussed during section 4.3. Finally, note that Saclay mocks have a better reproduction of the cross-correlation at small scales than LyaCoLoRe mocks which is clearly seen at the $0.50<\mu<$ 0.5 range, which is expected due to their QSO position sampling method differences discussed in section 2.1.
@@ 4.3 Astrophysical Contaminants
The DLA finder algorithm has been found to have purity and efficiency above the 90% level for systems detected in eBOSS DR16 data with $\log N_{HI}>20.1\ \rm{cm}^{-2}$ and redshift $z>2.2$ on quasars with with high mean flux $\bar{f_{\lambda}}>2\times 10^{-19}\,\mathrm{Wm^{-2}nm^{-1}}$ that generally have high SNR [52]. However, the performance on DESI data is still under investigation. Therefore, for the purposes of comparing the mock input DLA distribution with the distribution of detected DLAs in the EDR+M2 data, we have conservatively restricted our samples to those systems that fulfill $z_{\rm{DLA}}<z_{\rm{QSO}}$, $z_{\rm{DLA}}>2.6$, $\log N_{HI}>20.5\ \rm{cm}^{-2}$. We have also restricted the redshift of the host QSOs to $3.62.6<z_{\rm{QSO}}<$ 3.6 to match the minimum value of the DLA systems in the observed catalog and the maximum value of the Saclay mocks QSOs. Additionally, on data we restrict to DLA systems detected with a confidence larger than 50% by both the Convolutional Neural Network (CNN) and Gaussian Process (GP) methods of the DLA finder as was done in  [66]. All the mentioned restrictions applied on observed data yield a 34,053 QSOs sample and 2061 detected DLAs. In the case of our mocks we obtain a sample of approximately 35k QSOs and 2,100 DLAs with little mock-to-mock variation.
@@
The official BAL catalog of the EDR+M2 dataset contains 22.8k BAL quasars with Absorption Index $\rm{AI}>0$ detected by a BAL finder algorithm [47, 48] when restricting the sample to QSOs within the redshift range $3.61.8<z<$ 3.6. For mocks we have used a 16% probability of a quasar having a BAL in its spectrum, which results in approximately 22.1k $\rm{AI}>0$ BAL quasars for each mock realization. Figure 11 shows Absorption (AI) and Balnicity (BI) indices distributions as obtained by the BAL finder algorithm on data and one mock realization when using the restrictions $\rm{AI}>0$ (left panel) and $\rm{BI}>0$ (right panel). We find agreement between the number of $\rm{AI}>0$ objects on data and mocks, however there is a discrepancy on the shape of the AI distribution, which is discussed in [48] and might be due to the differences between the signal to noise ratio of DESI EDR+M2 and the eBOSS DR14 data used to create the BAL templates, we leave the study of the effect of SNR on the distribution shape for future DESI releases with higher statistics than the EDR+M2 sample. While the shape of the BI distribution seems to be in agreement, we have more $\rm{BI}>0$ BAL quasars on mocks than on observations, which we also attribute to the differences between the dataset used to generate the BAL templates and the EDR+M2 data, although this does not affect our results since the analysis pipeline BAL masking criteria is simply $\rm{AI}>0$. Future realizations of DESI Lyman-$\alpha$ mocks might require an update of these templates.
@@
The effect of the aforementioned Si is most noticeable in the first two $r_{\perp}$ bins of the 3D flux correlation functions $\xi_{\rm{3D}}$, which comprise the region closest to the line of sight. In figure 13 we show the contributions of Si II(1190), Si II(1193), Si II(1260), and Si III(1207) to the shape of the Lyman-$\alpha$ cross and auto correlation functions in these $r_{\perp}$ bins of one realization of the EDR+M2 mocks with metal contamination. These appear as prominent bumps on the cross-correlation at the scale -59, -52, 104 and -21 Mpc/h, respectively, calculated using equation 2.2 at the effective redshift $z_{eff}=2.37$. These bumps also appear in the auto-correlation at the absolute value of these same scales. We have multiplied the correlation function by $r$ to clearly show these bumps, but given the low statics of EDR+M2 data these are not as evident as in the best-fit model, also shown, computed using the velocity bias $b_{\eta,m}$ obtained for these particular mocks. For comparison with the DESI EDR+M2 data, we also show the measured correlations and fits obtained in [64]. The apparent mismatch in the first $r_{\perp}$ bin can be attributed to the small statistics in that particular bin, and the fact that the fit is performed using the full correlations.
Finally, in figure 14 we present the obtained biases of our 10 LyaCoLoRe and 10 Saclay mocks contaminated only with metals. Even though the overall dispersion of the resulting biases of the included metals are enough to be statistically consistent with EDR+M2 observed data for both type of mocks, there is a difference on the mean value of the biases obtained for Si II(1190), Si III(1207) bias of Saclay mocks and Si II(1260) in LyaCoLoRe mocks which contribute to the differences on the correlation functions discussed in section 4.2. The results on the metal velocity biases along with the results obtained on the 1D correlation suggests a need for a re-calibration of the relative absorption coefficient of these metals. The possibility of re-tuning the metal coefficients or exploring a non-linear tuning method is left for future DESI data releases.

### v3

@@ 3.1 Survey demography
By default we simulate over the whole DESI footprint, however if a specific area is to be simulated, such as fractions of the DESI footprint or the footprint from other surveys, we use an input number density function tabulated into HEALpix pixels [72] including only those pixels that cover the footprint to be simulated. Quickquasars then uses this information and random samples the number of objects from the available skewers to match the observed number density. We used this sub-sampling method as a first approximation for EDR+M2 mocks only, even though in principle it could affect the Lyman-$\alpha$ – QSO cross and QSO auto-correlations at large scales for the HEALpix pixel size used. [^12] We use HEALpix pixel of nside=16 which corresponds to a $\sim 250$ Mpc/h scale at redshift z=2.3. This sub-sampling method differs from the one used for eBOSS mocks in the DR16 analysis validation [10] and therefore does not affect its results. A more suitable method will be studied for future releases of DESI mocks.
@@
We include two options to assign exposure times. If it is not expected that all the simulated quasars have the same exposure time, we use an exposure time probability distribution function computed for each HEALpix pixel of the footprint to randomly assign a multiple of 1000 seconds to the selected quasars. Otherwise, we assign all the quasars the same exposure time, 4000 seconds in the case of complete DESI survey mocks.
@@ 3.2.3 Variance smoothing
This smoothing step is not part of specsim, so we emulate this process in quickquasars as follows: We subtract the source electron shot noise contribution from the total variance estimated by specsim, explained in section 3.2.2. We smooth this source electron contribution with a Gaussian kernel using Fast Fourier Transforms (FFTs), where the input array is padded with boundary values to prevent periodicity from distorting values at the edges. Finally, this smoothed source contribution is added to the total variance. The standard deviation of the Gaussian kernel $\sigma$ is an input parameter in quickquasars, which we set to $10~{}\text{\AA}$ by default based on the DESI spectroscopic data processing pipeline.
@@ 3.2.4 Observing conditions
The atmosphere model presented in section 3.2.1 depends on the assumed observing conditions such as seeing, air mass, moon illumination fraction, moon altitude and moon separation from the tile being simulated. For Lyman-$\alpha$ mocks we use the DESI dark-program optimal conditions setup where the seeing is set to $1.1\ \rm{arcsec}$, air mass to 1.0, and we set the moon illumination fraction to 0 while the moon altitude and separation from the tile are set to $-60\ \rm{degrees}$ and $180\ \rm{degrees}$, respectively. While these conditions might not be particularly realistic, they correspond to how the effective exposure time is defined for DESI observations. In other words, a 1000 seconds exposure simulated at these conditions provides the expected signal-to-noise ratio that an actual observation under different conditions aims to achieve.
@@ 3.2.5 The resolution matrix
The DESI spectrograph resolution is encoded in our data products in the form of a "resolution matrix" that is saved in the output spectra files. In the observed data case, this matrix is a result of the spectroscopic extraction algorithm and encodes the dispersion along wavelength due to the spectrograph finite PSF and the CCD pixel size. More details can be found in section 4.5 of [75]. This band-diagonal matrix  [^15] The resolution matrix elements far from the diagonal are zero since it combines only flux values from neighboring wavelength. applied to a high resolution spectrum results in a spectrum at the resolution of the observations. For mocks, each row of the resolution matrix can be approximated as a Gaussian of parameter $\sigma=(\sigma_{LSF}^{2}+\Delta\lambda^{2}/12)^{1/2}$, where $\sigma_{LSF}=0.73$Å at 4000Å and $\Delta\lambda=0.8$Å is the output wavelength bin size.
@@ 4 DESI EDR+M2 Lyman-$\alpha$ mocks comparison with data
In this section, we present the DESI Lyman-$\alpha$ EDR+M2 mocks, a collection of 40 different mock datasets: 20 for LyaCoLoRe and 20 for Saclay raw mocks, of which 10 mocks contain only Lyman-$\alpha$ absorption and 10 additionally include HCDs, BALs and metals. The metals were extracted directly from the transmission files for LyaCoLoRe, and added by quickquasars for Saclay. We include absorption due to Lyman-$\beta$, Si II(1190), Si II(1193), Si II(1260) and Si III(1207).
To produce all of these mocks, we extracted the observed footprint, quasar number density, and exposure time distribution from quasar catalog used for the DESI EDR+M2 Lyman-$\alpha$ 3D correlation functions measurement [76]. This catalog combines the data of DESI Early Data Release (EDR), collected during the survey validation phase, and the first two months of the Main Survey (DESI-M2). The resulting observed data sample consists of 318k QSO targets within the $60<z<$ 6 redshift range. Next, we show a comparison of the results of these mocks to those of the observed data.
@@ 4.1 Demographics
For mocks we only consider QSOs with redshift $3.81.8<z<$ 3.8, which results in a sample of 141k targets including 107k Lyman-$\alpha$ QSOs at $z>2.1$. We ignore targets in the regions explored during the EDR phase that fall outside the main DESI footprint (far north and far south) since these regions are not considered by the raw mocks. This makes no impact on our comparison since these targets represent only 0.9% of the sample. The resulting footprint of the EDR+M2 dataset with the considerations previously mentioned is shown in figure 4. By construction all of the produced mocks follow the same footprint, QSO number density and redshift-magnitude distributions as the observed data with negligible variations between realizations.
As mentioned in section 3.1, our mocks follow by construction the redshift-magnitude distribution from figure 1 which is slightly different from that measured on the EDR+M2 sample. This is due the fact that the EDR sample includes fainter targets than those expected to be measured during the main DESI survey (e.g DESI-M2). [^16] During the survey validation phase an extension of the r-band magnitude limit was tested to study the redshift distribution and population of fainter objects. However, this does not affect the quality of our mocks since the amount of this type of target represents a small fraction of the sample as can be seen in the left panel of figure 5 at $M_{r}>23.1$.
@@
Figure 5 shows the observed EDR+M2 data redshift (right) and r-band magnitude (left) distributions, and the corresponding results of one mock realization of each LyaCoLoRe and Saclay mocks. The differences between mocks and data distributions are below the 1% level and we obtain negligible dispersion between mock realizations. This validates the appropriateness of our methods of redshift sampling and random r-band magnitude assignation, described in section 3.1. We see a cutoff in the redshift distribution of the mocks at $z>3.8$ for LyaCoLoRe and $z>3.6$ for Saclay. This comes from the fact that we reach the redshift limit of the raw mocks at this point. The effect of these limits in our analysis and validations are negligible, since the QSOs at these redshifts represent roughly 1% of the observed sample.
To emulate the signal to noise ratio (SNR) of observed data, we assign random exposure times to the simulated quasars by using am exposure time distribution function, as explained in section 3.1. We obtained the distribution function used for our EDR+M2 mocks from the effective exposure time of observed data defined by spectroscopic data processing pipeline as $T_{\rm{eff}}=12.15\ \rm{seconds}\times\text{TSNR}^{2}_{\rm{LRG}}$, where $\text{TSNR}_{\rm{LRG}}$ is the LRG template signal to noise ratio (see section 4.14 in [75], for further details). Additionally, we use the measured throughput model achieved during the EDR phase (blue line in figure 3), which includes a dip feature at $\lambda\sim 440$ nm that affects the noise of the simulated spectra [74]. We also consider the effect of galactic extinction on spectra by adding an O’Donnell extinction model [77]. This in principle modifies the magnitude of the mock QSOs from the randomly assigned ones, and therefore modifies the resulting magnitude distribution. We address this issue by re-scaling all fluxes by a factor $F=f_{0}/f_{\rm{EBV}}$, where $f_{0}$ is the flux randomly assigned by our method, and $f_{\rm{EBV}}$ is the flux after galactic extinction has been applied.
Figure 6 shows the median SNR of a 75k randomly selected QSO sample in the Lyman-$\alpha$ forest region of observed data and a mock realization of both LyaCoLoRe and Saclay mocks. We have considered r-band magnitude and redshift bins of 0.5 and 0.1 width, respectively. We see a significant difference in the SNR of our mocks compared to data at bright low redshifts quasars ($z<2.2$, $M_{r}<22$), where we obtain higher SNR in our mocks than in observed data, which might be due to an underestimation in our instrumental noise model for bright low redshift quasars. We expect to study and improve on this issue for future mocks.
@@ 4.2 Correlation functions
The variance $\sigma_{q}$ for the auto-correlation has contributions from instrumental noise, $\sigma_{\rm{pip}}$, and the intrinsic variance of the Lyman-$\alpha$ forest flux transmission fluctuations, $\sigma_{\rm{LSS}}$, through
@@
where $\tilde{\sigma}_{\rm{pip},q}=\sigma_{\rm{pip},q}(\lambda)/\bar{F}C_{q}(\lambda)$, and $\eta(\lambda)$ is a correction factor to account for inaccuracies in the estimation of $\sigma_{\rm{pip}}$. We fix $\eta(\lambda)=1$ in our analysis following [78]. The $\sigma_{\rm{LSS}}$ parameter, which quantifies the noise introduced to our analysis by the small-scale fluctuations in the Lyman-$\alpha$ forest, is fitted iteratively at the same time as the rest-frame continuum $\bar{C}(\lambda_{\rm{RF}})$ and the quasar spectral diversity parameters $a_{q}$ and $b_{q}$ [78].
@@
Figure 7 shows the comparison of the distributions of the quasar spectral diversity parameters $a_{q}$ and $b_{q}/a_{q}$ as measured from data and mocks. We have restricted to those objects whose amplitude parameter is $a_{q}>0$. The difference between the $b_{q}/a_{q}$ distributions of data and mocks suggests that an update in the simqso broken power law model slopes and dispersions might be required to better reflect the quasar spectral diversity observed in the DESI Lyman-$\alpha$ sample.
@@
where $\gamma_{\rm{Ly\alpha}}=2.9$ [80], $\gamma_{\rm{QSO}}=1.44$ [81] and $\sigma_{\rm{mod}}^{2}$ is an extra parameter introduced to modulate the contribution of the variance of the Lyman-$\alpha$ transmission fluctuations $\sigma_{\rm{LSS}}$. We have fixed the value $\sigma_{\rm{mod}}^{2}=7.5$ as this value was found to optimize the precision on the results of the correlation functions for the EDR+M2 dataset [78].
@@
Figure 9 shows the Lyman-$\alpha$ auto (top) and Lyman-$\alpha$ – QSO cross (bottom) correlation functions of the measurements presented by [76] in four ranges of $\mu=r_{\parallel}/r$. We also show individual correlations of the 10 mock realizations with contaminants for both LyaCoLoRe and Saclay mocks and their median value. We have multiplied the correlations by $r^{2}$ for better appreciation of the BAO peak. Overall, the computed correlation functions of Saclay and LyaCoLoRe mocks are consistent with observations, however there are some clear differences between mocks and observed data, which we discuss next.
To begin with, we see significant differences at small scales specially at the $0.50<\mu<$ 0.5 range of the auto-correlation, and the $0.80.5<\mu<$ 0.8 and $0.950.8<\mu<$ 0.95 ranges of the cross-correlation, which may be attributed to different reasons. First, the Lyman-$\alpha$ and quasar biases measured on the EDR+M2 observed Lyman-$\alpha$ data, and the target values used to produce the raw mocks are different. The LyaCoLoRe mocks target the results of BOSS DR12, while Saclay mocks target eBOSS DR16. Second, as discussed in section 2.3, systems that go undetected by the DLA finder algorithm, and therefore are not masked, have a significant impact on the shape of the correlation functions, this is not the case for our mocks where we mask all of these features. Third, statistical errors in the measurement of quasar redshifts introduce spurious correlations that mostly affect the shape of the correlation functions in the $0.950.8<\mu<$ 0.95 and $10.95<\mu<$ 1 ranges for both auto and cross correlations [49], this effect was not included in our mocks. Fourth, instrumental noise has contributions to the shape of the correlation functions more predominantly at small scales of the $0.50<\mu<$ 0.5 range. In this regard, there exists the possibility of the model used to introduce instrumental noise not being fully representative of what is measured on data. These possible sources will be studied with mocks of future DESI releases with a higher signal-to-noise ratio. The differences at the $1.00.95<\mu<$ 1.0 range at approximately 20 and 60 Mpc/h in the auto-correlation and 110 Mpc/h in the cross-correlation might have some contributions by the value chosen to include metals in our mocks, this will be further discussed during section 4.3. Finally, note that Saclay mocks have a better reproduction of the cross-correlation at small scales than LyaCoLoRe mocks which is clearly seen at the $0.50<\mu<$ 0.5 range, which is expected due to their QSO position sampling method differences discussed in section 2.1.
@@ 4.3 Astrophysical Contaminants
The DLA finder algorithm has been found to have purity and efficiency above the 90% level for systems detected in eBOSS DR16 data with $\log N_{HI}>20.1\ \rm{cm}^{-2}$ and redshift $z>2.2$ on quasars with high mean flux $\bar{f_{\lambda}}>2\times 10^{-19}\,\mathrm{Wm^{-2}nm^{-1}}$ that generally have high SNR [64]. However, the performance on DESI data is still under investigation. Therefore, for the purposes of comparing the mock input DLA distribution with the distribution of detected DLAs in the EDR+M2 data, we have conservatively restricted our samples to those systems that fulfill $z_{\rm{DLA}}<z_{\rm{QSO}}$, $z_{\rm{DLA}}>2.6$, $\log N_{HI}>20.5\ \rm{cm}^{-2}$. We have also restricted the redshift of the host QSOs to $3.62.6<z_{\rm{QSO}}<$ 3.6 to match the minimum value of the DLA systems in the observed catalog and the maximum value of the Saclay mocks QSOs. Additionally, on data we restrict to DLA systems detected with a confidence larger than 50% by both the Convolutional Neural Network (CNN) and Gaussian Process (GP) methods of the DLA finder as was done in  [78]. All the mentioned restrictions applied on observed data yield a 34,053 QSOs sample and 2061 detected DLAs. In the case of our mocks we obtain a sample of approximately 35k QSOs and 2,100 DLAs with little mock-to-mock variation.
@@
The official BAL catalog of the EDR+M2 dataset contains 22.8k BAL quasars with Absorption Index $\rm{AI}>0$ detected by a BAL finder algorithm [59, 60] when restricting the sample to QSOs within the redshift range $3.61.8<z<$ 3.6. For mocks we have used a 16% probability of a quasar having a BAL in its spectrum, which results in approximately 22.1k $\rm{AI}>0$ BAL quasars for each mock realization. Figure 11 shows Absorption (AI) and Balnicity (BI) indices distributions as obtained by the BAL finder algorithm on data and one mock realization when using the restrictions $\rm{AI}>0$ (left panel) and $\rm{BI}>0$ (right panel). We find agreement between the number of $\rm{AI}>0$ objects on data and mocks, however there is a discrepancy on the shape of the AI distribution, this is discussed in [60] and might be due to the differences between the signal-to-noise ratio of DESI EDR+M2 and the eBOSS DR14 data used to create the BAL templates, we leave the study of the effect of SNR on the distribution shape for future DESI releases with higher statistics than the EDR+M2 sample. While the shape of the BI distribution seems to be in agreement, we have more $\rm{BI}>0$ BAL quasars on mocks than on observations, which we also attribute to the differences between the dataset used to generate the BAL templates and the EDR+M2 data, although this does not affect our results since the analysis pipeline BAL masking criteria is simply $\rm{AI}>0$. Future realizations of DESI Lyman-$\alpha$ mocks might require an update of these templates.
@@
The effect of the aforementioned metals is most noticeable in the first two $r_{\perp}$ bins of the 3D flux correlation functions $\xi_{\rm{3D}}$, which comprise the region closest to the line of sight. In figure 13 we show the contributions of Si II(1190), Si II(1193), Si II(1260), and Si III(1207) to the shape of the Lyman-$\alpha$ cross and auto correlation functions in these $r_{\perp}$ bins of one realization of the EDR+M2 mocks with metal contamination. These appear as prominent bumps on the cross-correlation at the scale -59, -52, 104 and -21 Mpc/h, respectively, calculated using equation 2.2 at the effective redshift $z_{eff}=2.37$. These bumps also appear in the auto-correlation at the absolute value of these same scales. We have multiplied the correlation function by $r$ to clearly show these bumps, but given the low statistics of the EDR+M2 data sample these are not as evident as in the best-fit model, also shown, computed using the velocity bias $b_{\eta,m}$ obtained for these particular mocks. For comparison with the DESI EDR+M2 data, we also show the measured correlations and fits obtained in [76]. The apparent mismatch in the first $r_{\perp}$ bin can be attributed to the small statistics in that particular bin, and the fact that the fit is performed using the full correlations.
Finally, in figure 14 we present the biases obtained from our 10 LyaCoLoRe and 10 Saclay mocks contaminated only with metals. Although the overall dispersion of the resulting biases of the included metals is enough to be statistically consistent with the observed data from EDR + M2 for both types of mocks, there is a difference in the mean value of the biases obtained for Si II(1190), Si III(1207) bias of Saclay mocks and Si II(1260) in LyaCoLoRe mocks which contribute to the differences on the correlation functions discussed in section 4.2. The results on the metal velocity biases along with the results obtained on the 1D correlation suggests a need for a re-calibration of the relative absorption coefficient of these metals. The possibility of re-tuning the metal coefficients or exploring a non-linear tuning method is left for future DESI data releases.